OpenStudy (nikki_skellington):

Alinn flipped a coin 20 times. The coin landed heads up 12 times and tails up 8 times.

OpenStudy (nikki_skellington):

What is the theoretical probability of the coin landing heads up? Show your work.

OpenStudy (supersmart1001):

ii dont know what people r doing but thats what they got

OpenStudy (nikki_skellington):

hmm...ok thx for helping though :D

OpenStudy (supersmart1001):

i can get someone here

OpenStudy (nikki_skellington):

ok

OpenStudy (supersmart1001):

@valadin1 here is a question for u care to answer it

OpenStudy (nikki_skellington):

hes offline

OpenStudy (supersmart1001):

no im going to tell him to get online

OpenStudy (nikki_skellington):

o..ok

OpenStudy (nikki_skellington):

?

OpenStudy (nikki_skellington):

@welshfella could u help?

OpenStudy (welshfella):

the experimental probability is the one derived from an experiment. Here its P(H) = 12/20 = 0.6 and for tails P(T) = 8/20 = 0.4

OpenStudy (welshfella):

As a fair coiun has 2 sides the probability in theory of H = 1/2 and of T = 1/2

OpenStudy (welshfella):

theoretical probabilty is 0.5 for H and T.

OpenStudy (nikki_skellington):

ok so is it 40% for (t) @welshfella

OpenStudy (welshfella):

that the experimental probability YES

OpenStudy (nikki_skellington):

so 40% for (t)...

OpenStudy (welshfella):

yep and 0.6 for heads

OpenStudy (nikki_skellington):

ok thx so much :)

OpenStudy (welshfella):

yw